# Deployable Reinforcement Learning with Variable Control Rate

### v1

@@ Introduction
Temporal aspects of reinforcement learning (RL), such as the duration of the execution of each action or the time needed for observations, are frequently overlooked. This oversight arises from the foundational hypothesis of the Markov Decision Process (MDP), which assumes the independence of each action undertaken by the agent (Norris 1998). As depicted in the top section of Figure 1, conventional RL primarily focuses on training an action policy, generally neglecting the intricacies of policy implementation. Some prior research approached the problem by splitting their control algorithm into two distinct components (Williams et al. 2017): a learning part responsible for proposing an action policy, and a control part responsible for implementing the policy (Yang et al. 2018; Zanon and Gros 2020; Mahmood et al. 2018).
Translating action policies composed of discrete time steps into real-world applications generally means using a fixed control rate (e.g., 10 Hz). Practitioners typically choose the control rate based on their experience and the specific needs of each application, often without considering adaptability or responsiveness to changing environmental conditions. Imagine driving a car in a straight line in an extensive environment without obstacles: very few control actions are needed. On the contrary, driving in tight spaces with low tolerances and following complex paths can require a higher control frequency.Setting the control frequency is set to a fixed value requires it to be the worst-case (or an average is safety is not a concern) scenario that allows controllability of the system. In practical applications of reinforcement learning, especially in scenarios with constrained onboard computer resources, maintaining a consistently high fixed control rate can limit the availability of computing resources for other tasks and significantly increase energy consumption.
@@
The reward function is defined as: $R=\alpha_{t}\cdot R_{t}-\alpha_{\varepsilon}\cdot R_{\varepsilon}-\alpha_{\tau}\cdot R_{\tau}$ (1) where $R_{t}$ = $r$, r is the value determined by task setting; $R_{\varepsilon}=\varepsilon$, $\varepsilon$ is the energy cost of one time step; $R_{\tau}=\tau$, $\tau$ is the time taken to execute a time step. $\alpha_{t},\alpha_{\varepsilon},\alpha_{\tau}$ are parametric weighting factors. We determine the optimal policy $\pi^{*}$, which maximizes the reward $R$.
@@ Appendix A Appendix:A Validation Environment Details
State We have only one starting point, endpoint, and obstacle in the set environment. The positions of the endpoint and obstacle are randomized in each episode. $S_{t}=(Pos,Obs,Goal,Speed,Time,Force)$ (7) Where: $Pos$ = Position Data of The Agent in X and Y Direction, $Obs$ = Position Data of The Obstacle in X and Y Direction, $Goal$ = Position Data of The Goal in X and Y Direction, $Speed$ = Spped Data of The Agent for Current Step, $Time$ = The Duration data of The Agent to Execute The Last Time Step, $Force$ = Force Data of The Agent for Last Time Step in X and Y Direction.
@@
Action Each action should have a corresponding execution duration. $a_{t}=(a_{t},a_{fx},a_{fy})$ (8) Where: $a_{t}$ = The time of The Agent to implement current action, $a_{fx}$ = The Force of The Agent in X Coordinate, $a_{fy}$ = The Force of The Agent in Y Coordinate.

### v2

@@ Introduction
Temporal aspects of reinforcement learning (RL), such as the duration of the execution of each action or the time needed for observations, are frequently overlooked. This oversight arises from the foundational hypothesis of the Markov Decision Process (MDP), which assumes the independence of each action undertaken by the agent (Norris 1998). As depicted in the top section of Figure 1, conventional RL primarily focuses on training an action policy, generally neglecting the intricacies of policy implementation. Some prior research approached the problem by splitting their control algorithm into two distinct components (Williams et al. 2017): a learning part responsible for proposing an action policy, and a control part responsible for implementing the policy (Yang et al. 2018; Zanon and Gros 2020; Mahmood et al. 2018).  [^1] Our paper is presented at Deployable AI Workshop at AAAI-2024
Translating action policies composed of discrete time steps into real-world applications generally means using a fixed control rate (e.g., 10 Hz). Practitioners typically choose the control rate based on their experience and the specific needs of each application, often without considering adaptability or responsiveness to changing environmental conditions. Imagine driving a car in a straight line in an extensive environment without obstacles: very few control actions are needed. On the contrary, driving in tight spaces with low tolerances and following complex paths can require a higher control frequency. Setting the control frequency is set to a fixed value requires it to be the worst-case (or an average if safety is not a concern) scenario that allows controllability of the system. In practical applications of reinforcement learning, especially in scenarios with constrained onboard computer resources, maintaining a consistently high fixed control rate can limit the availability of computing resources for other tasks and significantly increase energy consumption.
@@
The reward function is defined as: $R=\alpha_{t}\cdot R_{t}-\alpha_{\varepsilon}\cdot R_{\varepsilon}-\alpha_{\tau}\cdot R_{\tau}$ (1) where
$R_{t}$ = $r$, r is the value determined by task setting;
$R_{\varepsilon}=\varepsilon$, $\varepsilon$ is the energy cost of one time step;
$R_{\tau}=\tau$, $\tau$ is the time taken to execute a time step.
$\alpha_{t},\alpha_{\varepsilon},\alpha_{\tau}$ are parametric weighting factors.
We determine the optimal policy $\pi^{*}$, which maximizes the reward $R$.
@@ Appendix A Appendix:A Validation Environment Details
State We have only one starting point, endpoint, and obstacle in the set environment. The positions of the endpoint and obstacle are randomized in each episode.
$S_{t}=(Pos,Obs,Goal,Speed,Time,Force)$ (7)
Where:
$Pos$ = Position Data of The Agent in X and Y Direction,
$Obs$ = Position Data of The Obstacle in X and Y Direction,
$Goal$ = Position Data of The Goal in X and Y Direction,
$Speed$ = Spped Data of The Agent for Current Step,
$Time$ = The Duration data of The Agent to Execute The Last Time Step,
$Force$ = Force Data of The Agent for Last Time Step in X and Y Direction.
@@
Action Each action should have a corresponding execution duration.
$a_{t}=(a_{t},a_{fx},a_{fy})$ (8)
Where:
$a_{t}$ = The time of The Agent to implement current action,
$a_{fx}$ = The Force of The Agent in X Coordinate,
$a_{fy}$ = The Force of The Agent in Y Coordinate.